Divide $6.00 among 3 students so that when Amy gets 3 cents, Bono gets 5 cents, and Teri gets 7 cents. how much does each receive?
Answer by stanbon(75887)

Divide $6.00 among 3 students so that when Amy gets 3 cents, Bono gets 5 cents, and Teri gets 7 cents. how much does each receive?
--------------------------------
Amy: 3x
Bono: 5x
Teri: 7x
--------------------------------
Equation:
3x + 5x + 7x = 600
15x = 600
x = 40 cents
--------------------
Amy: 3x = $1.20
Bono:5x = 5*40 = $2.00
Teri:7x = 7*40 = $2.80
===============================
Cheers,
Stan H.